OpenStudy (candyme):

Four ninths of a number c increased by 4 is 18 less than one eighth times another number d. Solve for c. A. c=9/32 d +31 1/2 B. c=4/72 d + 4/72 C. c=9/32 d - 49 1/2 D. c=4/72 d - 31 1/2 I would like to see HOW the answer was reached, not just the answer.

OpenStudy (anonymous):

4/9 c +4 = 1/8 d -18 (4/9)c = (1/8)d -22 c = (9/32)d -(99/2) c= (9/32)d - 49 1/2 The answer is choice C.

OpenStudy (candyme):

Sorry, I'm still a bit confused. Can you explain what you did in steps 2-3?

OpenStudy (anonymous):

When you move +4 to the right side of the equation, it is -4. So, -4-18= -22 In step 3, it is c=(9/4)*{(1/8)d -22} Multiply and you get the expression in step c

OpenStudy (candyme):

Oh! Now I see what happened! Thank you soo much! :)